# Mask-informed Deep Contrastive Incomplete Multi-view Clustering

### v1

@@ Abstract
Multi-view clustering (MvC) utilizes information from multiple views to uncover the underlying structures of data. Despite significant advancements in MvC, mitigating the impact of missing samples in specific views on the integration of knowledge from different views remains a critical challenge. This paper proposes a novel Mask-informed Deep Contrastive Incomplete Multi-view Clustering (Mask-IMvC) method, which elegantly identifies a view-common representation for clustering. Specifically, we introduce a mask-informed fusion network that aggregates incomplete multi-view information while considering the observation status of samples across various views as a mask, thereby reducing the adverse effects of missing values. Additionally, we design a prior knowledge-assisted contrastive learning loss that boosts the representation capability of the aggregated view-common representation by injecting neighborhood information of samples from different views. Finally, extensive experiments are conducted to demonstrate the superiority of the proposed Mask-IMvC method over state-of-the-art approaches across multiple MvC datasets, both in complete and incomplete scenarios.
Multi-view clustering, incomplete multi-view clustering
[…]
1 Introduction
@@
Multi-view data commonly appears in many real-world applications with the rapid development of diverse feature descriptors and data collectors (Zhou et al., 2024; Wang, 2021; Wang et al., 2023b). For instance, the visual, text, and audio information are extracted and exploited in various video understanding tasks (Li et al., 2024b; Schiappa et al., 2023). To handle this type of data, multi-view learning methods are proposed to integrate the information from various views to achieve compatible performance than the approaches individually trained on a single view (Li et al., 2024b). Multi-view clustering (MvC), one of the kinds of multi-view learning methods, groups the data points into their respective clusters by exploring the comprehensive information among various views without labels, and has received much attention in recent years (Zhou et al., 2024; Zou et al., 2023, 2024).
@@
In many real-world applications, the samples in particular views may be absent due to hardware faults or interference signals during information collection, resulting in incomplete multi-view data (Wen et al., 2024). The incomplete multi-view clustering (IMvC) methods are introduced to address this type of data and have witnessed great progress recently (Wen et al., 2024; Lin et al., 2021). Previous IMvC approaches mainly perform cross-view imputation to recover the absent information among different views based on cross-view semantic consistency as illustrated in Fig. 1 (a). For example, Li et al. (Li et al., 2023) introduced a prototype-based imputation manner that utilizes a dual-attention equipped with a contrastive loss to seek prototypes of each view for recovering the missing views. Pu et al. (Pu et al., 2024) designed an adaptive imputation layer that leverages the soft cluster assignments across different views and the global cluster centroids to impute the missing views in a latent embedding space. However, the imputation procedure of missing views without the guidance of true data distributions inevitably introduces noise information and consequentially reduces the quality of imputed views. As a result, the low-quality imputed views further degrade the learning process of the IMvC methods, leading to an under-optimal view-common representation for clustering.
@@
Different from previous methods, we propose to cluster the incomplete multi-view data without the missing view imputation to avoid introducing inaccurate information. To achieve this goal, the following challenge needs to be addressed: How can we reduce the impact of missing values across different views when merging information from multiple sources while learning a discriminative view-common feature representation for data clustering? The missing views are difficult to be directly fused duo the missing values. Even some imputation manners are exploited to handle the missing views so that the multi-view features are easier to be aggregated, the uncontrollable information underlying the missing views contains inevitable adverse effects during the fusion process. Thus, it is emerging to design a flexible and effective multi-view fusion strategy to learn and integrate multi-view information when facing either complete or incomplete parts of the data. The final target of multi-view information fusion is to seek a view-common representation for clustering. The prior knowledge, e.g. the manifold structures (Meilă & Zhang, 2024) of different views, is valuable for representation learning since the prior knowledge helps to regularize the aggregated view-common feature representation with a compact and discriminative structure that polishes the noise information. Yet previous methods usually ignore utilizing such prior structural knowledge to boost the aggregated view-common features in the IMvC task.
@@
Based on the above discussion, we propose Mask-IMvC (short for Mask-informed Deep Contrastive Incomplete Multi-view Clustering), a novel solution that flexibly and effectively integrates the information among incomplete multi-view data for clustering. As shown in Fig 1 (b), we first introduce a mask-informed fusion network that takes the observation status of samples across various views into consideration to formulate a mask, removing the contributions missing sample during the multi-view information aggregation. Afterward, a prior knowledge-assisted contrastive learning is designed, in which the neighbor correlations of each sample captured from different views are injected into the learned view-common representation with a re-weighted contrastive loss to achieve a more compact and clear cluster structure. As s result, the incomplete multi-view information is elegantly aggregated into a view-common feature representation for clustering. The contributions are summarized as follows,
@@
2 Related Work
2.1 Incomplete Multi-view Clustering
Existing IMvC approaches can be generally grouped into: traditional IMvC methods (Wen et al., 2020a; Yin & Sun, 2021; Wang et al., 2022) and deep IMvC algorithms (Yang et al., 2022; Xu et al., 2024a; Yan et al., 2024).
Traditional IMvC methods commonly build upon matrix factorization, kernel learning, and graph learning technologies. Matrix factorization-based approaches derive a view-shared representation across diverse views via the matrix factorization techniques (Hu & Chen, 2019; Wen et al., 2024; Khan et al., 2024). For instance, an adaptive feature weighting manner is inserted into matrix factorization to mitigate the effects of redundant and noisy features on the learned view-shared representation, which is also regularized with a graph-embedded consensus constraint to preserve the structural information inherent in incomplete multi-view data. Kernel-based methods (Liu, 2024; Li et al., 2024a) utilizes a set of per-computed kernels to measure diverse views, tending to formulate a unified kernel through linear or non-linear combinations of predefined kernels to capture the clustering results. Liu et al. (Liu, 2024) jointly conduct incomplete kernel matrices imputation and alignment to capture an advanced clustering representation. Graph-based methods (Yang et al., 2024; Du et al., 2024; Li et al., 2022) integrate graph similarities captured from various views via either self-representation (Elhamifar & Vidal, 2013) or adaptive neighbor graph learning (Nie et al., 2014) manner, ultimately obtaining the clustering results through spectral clustering. Li et al. (Li et al., 2022) explored the cross-view information to refine the graph structure by leveraging the tensor nuclear norm. Deep learning-based methods (Wen et al., 2020b; Xu et al., 2022a; Li et al., 2023) leverage the powerful representation capabilities of deep neural networks to derive consensus clustering results from multi-view data. For example, Lin et al. (Lin et al., 2021) learned the consensus representation across diverse views by contrastive learning and recovered the missing views by cross-view prediction. Xue et al. (Xue et al., 2024) developed a multi-graph contrastive regularization to reduce abundant correlations across multiple views and learn discriminative representations for clustering.
@@ 2.1 Incomplete Multi-view Clustering
2.2 Contrastive Learning
Contrastive learning is a novel self-supervised learning paradigm that has achieved significant success across various computer vision and machine learning tasks (Krishnan et al., 2022; Liu et al., 2021b). Its core principle involves pushing samples away from their negative anchors while pulling samples closer to their positive anchors, thereby creating a discriminative representation for downstream tasks (Khosla et al., 2020). Over the past few years, different contrastive learning approaches have emerged, including MoCo (He et al., 2020), SimCLR (Chen et al., 2020), and SwAV (Caron et al., 2020). For a comprehensive overview of additional methods, we refer to the survey in (Gui et al., 2024).
Recently, inspired by the robust feature learning capabilities of self-supervised learning, contrastive loss has been extensively applied in MvC. For instance, Xu et al.(Xu et al., 2022b) aligned multi-view information from both high-level semantics and low-level features through contrastive learning, effectively capturing common semantics for clustering. Additionally, the work in(Trosten et al., 2023) explored the effectiveness of self-supervision and contrastive alignment within the multi-view clustering task. Luo et al. (Luo et al., 2024) first fused multi-view information at the data level and then applied data augmentation for the fused data, making a shared feature extractor with a simple contrastive learning paradigm to capture robust features for clustering. However, previous methods typically treat the same samples from different views as positive pairs, while diverse samples across multiple views are considered negative pairs. This manner lacks the flexibility to leverage sample connection probabilities to enhance feature representation learning. Thus, we propose prior knowledge-assisted contrastive learning to inject the sample connection probabilities from diverse views into the view-common feature representation with a carefully designed re-weighted contrastive loss. In such a manner, the view-common feature representation full of structural information is beneficial for recovering the underlying cluster structure of data.
@@ 3.1 Notation
To transfer the observed parts $\{\bar{\mathbf{Z}}^{v}\}_{v=1}^{V}$ and view missing parts $\{\hat{\mathbf{Z}}^{v}\}_{v=1}^{V}$ back into the original sample orders in $\{\mathbf{X}^{v}\}_{v=1}^{V}$, we formulate two kinds of transition matrices, i.e, $\{\bar{\mathbf{T}}^{v}\in\mathbb{R}^{N_{o}^{v}\times N}\}_{v=1}^{V}$ and $\{\hat{\mathbf{T}}^{v}\in\mathbb{R}^{N_{m}^{v}\times N}\}_{v=1}^{V}$ as,
$\bar{t}_{ij}^{v}=\begin{cases}1&\text{if }\bar{\mathbf{z}}_{i}^{v}\text{ is inthe position of }\mathbf{x}_{j}^{v},\\ 0&\text{otherwise},\end{cases}$ (2)
$\hat{t}_{ij}^{v}=\begin{cases}1&\text{if }\hat{\mathbf{z}}_{i}^{v}\text{ is inthe position of }\mathbf{x}_{j}^{v},\\ 0&\text{otherwise},\end{cases}$ (3)
where $\bar{t}_{ij}^{v}$ and $\hat{t}_{ij}^{v}$ are the $ij$-th elements in $\bar{\mathbf{T}}^{v}$ and $\hat{\mathbf{T}}^{v}$, respectively. Based on them, we can permute the observed parts $\{\bar{\mathbf{Z}}^{v}\}_{v=1}^{V}$ and view missing parts $\{\hat{\mathbf{Z}}^{v}\}_{v=1}^{V}$ back into the original feature orders as,
$\mathbf{X}^{v}=\bar{\mathbf{Z}}^{v}\bar{\mathbf{T}}^{v}+\hat{\mathbf{Z}}^{v}\hat{\mathbf{T}}^{v}$ (4)
3.2 Overview
@@
View-specific encoders: The graph convolutional networks (GCNs) are widely used in many data mining tasks, aggregating the neighbor information to learn features (Xia et al., 2021). In our method, we first construct $k$-nearest neighbor graphs $\{\mathbf{A}^{v}\in\mathbb{R}^{N_{o}^{v}\times N_{o}^{v}}\}_{v=1}^{V}$ from the observed parts $\{\bar{\mathbf{Z}}\}_{v=1}^{V}$ by the adaptive neighbor learning method (Nie et al., 2014). The view-specific encoders $\{f^{v}(\bar{\mathbf{Z}}^{v},\mathbf{A}^{v};\Theta^{v})\}_{v=1}^{V}$, with $\Theta^{v}$ denoting the parameters, are implemented by several GCN layers to merge the neighbor information of each sample to learn view-specific latent representations. One of the GCN layers is formulated as follows,
@@
Mask-informed fusion: After the view-specific feature extraction, the obtained multiple features are permuted as the same orders as the original features as by transition matrices $\{\bar{\mathbf{T}}^{v}\in\mathbb{R}^{N_{o}^{v}\times N}\}_{v=1}^{V}$ as,
$\bar{\mathbf{H}}^{v}=\mathbf{H}^{v}\bar{\mathbf{T}}^{v}.$ (6)
@@
where $\mathbf{f}_{i}$ is the $i$-th samples of view-common features $\mathbf{F}$. In such an above manner, incomplete information is automatically filtered to avoid adverse impacts on the view-common features. To further boost the representation capability of $\mathbf{F}$, we introduced a prior knowledge-assisted contrastive learning manner that is detailed depicted in Section 3.3.
View-specific decoders: The view-specific decoders are introduced to reconstruct the observed parts in the original IMvC data. To this end, we first obtain the observed parts of each view in the view-common feature $\mathbf{F}$ by transition matrices $\{\bar{\mathbf{T}}^{v}\}_{v=1}^{V}$ as,
$\bar{\mathbf{F}}^{v}=\bar{\mathbf{T}}^{v\top}\mathbf{F},$ (8)
@@
3.3 Prior Knowledge-assisted Contrastive Learning
The similarities among samples defined from the original features are critical prior knowledge for feature representation learning (He & Niyogi, 2003). In this section, we inject the prior knowledge into the view-common feature representation with useful self-supervised contrastive learning to further boost its representation capability. To this end, we formulate three key components, i.e., the clustering projection layer, mask-informed graph fusion, and contrastive loss.
Clustering projection layer: In this part, we first transfer the view-common representation into the cluster space with a single linear layer. Then, the QR decomposition (Gander, 1980) is introduced to guarantee the orthogonal property of the clustering indicator matrix $\mathbf{Y}\in\mathbb{R}^{N\times C}$, where $C$ is the number of clusters. And the clustering distributions of the incomplete multi-view data can be revealed in $\mathbf{Y}$.
Mask-informed graph fusion: To flexibly merge multiple similarity graphs constructed from IMvC data, view-specific graphs $\{\mathbf{A}^{v}\}_{v=1}^{V}$ are first transferred based on transition matrices $\{\bar{\mathbf{T}}^{v}\}_{v=1}^{V}$ as,

### v2

@@ Abstract
Multi-view clustering (MvC) utilizes information from multiple views to uncover the underlying structures of data. Despite significant advancements in MvC, mitigating the impact of missing samples in specific views on the integration of knowledge from different views remains a critical challenge. This paper proposes a novel Mask-informed Deep Contrastive Incomplete Multi-view Clustering (Mask-IMvC) method, which elegantly identifies a view-common representation for clustering. Specifically, we introduce a mask-informed fusion network that aggregates incomplete multi-view information while considering the observation status of samples across various views as a mask, thereby reducing the adverse effects of missing values. Additionally, we design a prior knowledge-assisted contrastive learning loss that boosts the representation capability of the aggregated view-common representation by injecting neighborhood information of samples from different views. Finally, extensive experiments are conducted to demonstrate the superiority of the proposed Mask-IMvC method over state-of-the-art approaches across multiple MvC datasets, both in complete and incomplete scenarios. The demo code for our work will be publicly available at https://github.com/guanyuezhen/Mask-IMvC.
Index Terms:
Multi-view clustering, Incomplete multi-view clustering, Prior knowledge mining, Contrastive learning.
I Introduction
@@
Multi-view data commonly appears in many real-world applications with the rapid development of diverse feature descriptors and data collectors [1, 2, 3]. For instance, the visual, text, and audio information are extracted and exploited in various video understanding tasks [4, 5]. To handle this type of data, multi-view learning methods are proposed to integrate the information from various views to achieve compatible performance than the approaches individually trained on a single view [4]. Multi-view clustering (MvC), one of the kinds of multi-view learning methods, groups the data points into their respective clusters by exploring the comprehensive information among various views without labels, and has received much attention in recent years [6, 7, 8].
@@
In many real-world applications, the samples in particular views may be absent due to hardware faults or interference signals during information collection, resulting in incomplete multi-view data [9]. The incomplete multi-view clustering (IMvC) methods are introduced to address this type of data and have witnessed great progress recently [9, 10]. Previous IMvC approaches mainly perform cross-view imputation to recover the absent information among different views based on cross-view semantic consistency as illustrated in Fig. 1 (a). For example, Li et al. [11] introduced a prototype-based imputation manner that utilizes a dual-attention equipped with a contrastive loss to seek prototypes of each view for recovering the missing views. Pu et al. [12] designed an adaptive imputation layer that leverages the soft cluster assignments across different views and the global cluster centroids to impute the missing views in a latent embedding space. However, the imputation procedure of missing views without the guidance of true data distributions inevitably introduces noise information and consequentially reduces the quality of imputed views. As a result, the low-quality imputed views further degrade the learning process of the IMvC methods, leading to an under-optimal view-common representation for clustering.
@@
Different from previous methods, we propose to cluster the incomplete multi-view data without the missing view imputation to avoid introducing inaccurate information. To achieve this goal, the following challenge needs to be addressed: How can we reduce the impact of missing values across different views when merging information from multiple sources while learning a discriminative view-common feature representation for data clustering? The missing views are difficult to be directly fused due the missing values. Even some imputation manners are exploited to handle the missing views so that the multi-view features are easier to be aggregated, the uncontrollable information underlying the missing views contains inevitable adverse effects during the fusion process. Thus, it is emerging to design a flexible and effective multi-view fusion strategy to learn and integrate multi-view information when facing either complete or incomplete parts of the data. The final target of multi-view information fusion is to seek a view-common representation for clustering. The prior knowledge, e.g. the manifold structures [13] of different views, is valuable for representation learning since the prior knowledge helps to regularize the aggregated view-common feature representation with a compact and discriminative structure that polishes the noise information. Yet previous methods usually ignore utilizing such prior structural knowledge to boost the aggregated view-common features in the IMvC task.
@@
Based on the above discussion, we propose Mask-IMvC (short for Mask-informed Deep Contrastive Incomplete Multi-view Clustering), a novel solution that flexibly and effectively integrates the information among incomplete multi-view data for clustering. As shown in Fig 1 (b), we first introduce a mask-informed fusion network that takes the observation status of samples across various views into consideration to formulate a mask, removing the contributions of missing sample during the multi-view information aggregation. Afterward, a prior knowledge-assisted contrastive learning is designed, in which the neighbor correlations of each sample captured from different views are injected into the learned view-common representation with a re-weighted contrastive loss to achieve a more compact and clear cluster structure. As s result, the incomplete multi-view information is elegantly aggregated into a view-common feature representation for clustering. The contributions are summarized as follows,
@@
II Related Work
II-A Incomplete Multi-view Clustering
Existing IMvC approaches can be generally grouped into: traditional IMvC methods [14, 15, 16] and deep IMvC algorithms [17, 18, 8].
Traditional IMvC methods commonly build upon matrix factorization, kernel learning, and graph learning technologies. Matrix factorization-based approaches derive a view-shared representation across diverse views via the matrix factorization techniques [19, 9, 20]. For instance, an adaptive feature weighting manner is inserted into matrix factorization to mitigate the effects of redundant and noisy features on the learned view-shared representation, which is also regularized with a graph-embedded consensus constraint to preserve the structural information inherent in incomplete multi-view data. Kernel-based methods [21, 22] utilizes a set of per-computed kernels to measure diverse views, tending to formulate a unified kernel through linear or non-linear combinations of predefined kernels to capture the clustering results. Liu et al. [21] jointly conduct incomplete kernel matrices imputation and alignment to capture an advanced clustering representation. Graph-based methods [23, 24, 25] integrate graph similarities captured from various views via either self-representation [26] or adaptive neighbor graph learning [27] manner, ultimately obtaining the clustering results through spectral clustering. Li et al. [25] explored the cross-view information to refine the graph structure by leveraging the tensor nuclear norm. Deep learning-based methods [28, 29, 11] leverage the powerful representation capabilities of deep neural networks to derive consensus clustering results from multi-view data. For example, Lin et al. [10] learned the consensus representation across diverse views by contrastive learning and recovered the missing views by cross-view prediction. Xue et al. [30] developed a multi-graph contrastive regularization to reduce abundant correlations across multiple views and learn discriminative representations for clustering.
@@ II-A Incomplete Multi-view Clustering
II-B Contrastive Learning
Contrastive learning is a novel self-supervised learning paradigm that has achieved significant success across various computer vision and machine learning tasks [31, 32]. Its core principle involves pushing samples away from their negative anchors while pulling samples closer to their positive anchors, thereby creating a discriminative representation for downstream tasks [33]. Over the past few years, different contrastive learning approaches have emerged, including MoCo [34], SimCLR [35], and SwAV [36]. For a comprehensive overview of additional methods, we refer to the survey in [37].
Recently, inspired by the robust feature learning capabilities of self-supervised learning, contrastive loss has been extensively applied in MvC. For instance, Xu et al.[38] aligned multi-view information from both high-level semantics and low-level features through contrastive learning, effectively capturing common semantics for clustering. Additionally, the work in[39] explored the effectiveness of self-supervision and contrastive alignment within the multi-view clustering task. Luo et al. [40] first fused multi-view information at the data level and then applied data augmentation for the fused data, making a shared feature extractor with a simple contrastive learning paradigm to capture robust features for clustering. However, previous methods typically treat the same samples from different views as positive pairs, while diverse samples across multiple views are considered negative pairs. This manner lacks the flexibility to leverage sample connection probabilities to enhance feature representation learning. Thus, we propose prior knowledge-assisted contrastive learning to inject the sample connection probabilities from diverse views into the view-common feature representation with a carefully designed re-weighted contrastive loss. In such a manner, the view-common feature representation full of structural information is beneficial for recovering the underlying cluster structure of data.
@@ III-A Notation
To transfer the observed parts $\{\bar{\mathbf{Z}}^{v}\}_{v=1}^{V}$ and view missing parts $\{\hat{\mathbf{Z}}^{v}\}_{v=1}^{V}$ back into the original sample orders in $\{\mathbf{X}^{v}\}_{v=1}^{V}$, we formulate two kinds of permutation matrices, i.e, $\{\bar{\mathbf{T}}^{v}\in\mathbb{R}^{N_{o}^{v}\times N}\}_{v=1}^{V}$ and $\{\hat{\mathbf{T}}^{v}\in\mathbb{R}^{N_{m}^{v}\times N}\}_{v=1}^{V}$, in which their elements are defined as,
$\bar{t}_{ij}^{v}=\begin{cases}1&\text{if }\bar{\mathbf{z}}_{i}^{v}\text{ is transferred to }\mathbf{x}_{j}^{v},\\ 0&\text{otherwise},\end{cases}$ (2)
$\hat{t}_{ij}^{v}=\begin{cases}1&\text{if }\hat{\mathbf{z}}_{i}^{v}\text{ is transferred to }\mathbf{x}_{j}^{v},\\ 0&\text{otherwise},\end{cases}$ (3)
where $\bar{t}_{ij}^{v}$ and $\hat{t}_{ij}^{v}$ are the $ij$-th elements in $\bar{\mathbf{T}}^{v}$ and $\hat{\mathbf{T}}^{v}$, respectively. $\bar{\mathbf{z}}_{i}^{v}$ and $\hat{\mathbf{z}}_{i}^{v}$ represent the $i$-th samples of $\bar{\mathbf{Z}}^{v}$ and $\hat{\mathbf{Z}}^{v}$, respectively. $\mathbf{x}_{j}^{v}$ denotes the $j$-th sample of $\mathbf{X}^{v}$. Then, we can permute the observed parts $\{\bar{\mathbf{Z}}^{v}\}_{v=1}^{V}$ and view missing parts $\{\hat{\mathbf{Z}}^{v}\}_{v=1}^{V}$ back into the original feature orders as,
$\mathbf{X}^{v}=\bar{\mathbf{T}}^{v\top}\bar{\mathbf{Z}}^{v}+\hat{\mathbf{T}}^{v\top}\hat{\mathbf{Z}}^{v}$ (4)
III-B Overview
@@
View-specific encoders: The graph convolutional networks (GCNs) are widely used in many data mining tasks, aggregating the neighbor information to learn features [41]. In our method, we first construct $k$-nearest neighbor graphs $\{\mathbf{A}^{v}\in\mathbb{R}^{N_{o}^{v}\times N_{o}^{v}}\}_{v=1}^{V}$ from the observed parts $\{\bar{\mathbf{Z}}\}_{v=1}^{V}$ by the adaptive neighbor learning method [27]. The view-specific encoders $\{f^{v}(\bar{\mathbf{Z}}^{v},\mathbf{A}^{v};\Theta^{v})\}_{v=1}^{V}$, with $\Theta^{v}$ denoting the parameters, are implemented by several GCN layers to merge the neighbor information of each sample to learn view-specific latent representations. One of the GCN layers is formulated as follows,
@@
Mask-informed fusion: After the view-specific feature extraction, the obtained multiple features are permuted as the same orders as the original features as by permutation matrices $\{\bar{\mathbf{T}}^{v}\}_{v=1}^{V}$ as,
$\bar{\mathbf{H}}^{v}=\bar{\mathbf{T}}^{v\top}\mathbf{H}^{v}.$ (6)
@@
where $\mathbf{f}_{i}$ is the $i$-th samples of view-common features $\mathbf{F}$. In such an above manner, incomplete information is automatically filtered to avoid adverse impacts on the view-common features. To further boost the representation capability of $\mathbf{F}$, we introduced a prior knowledge-assisted contrastive learning manner that is detailed depicted in Section III-C.
View-specific decoders: The view-specific decoders are introduced to reconstruct the observed parts in the original IMvC data. To this end, we first obtain the observed parts of each view in the view-common feature $\mathbf{F}$ by permutation matrices $\{\bar{\mathbf{T}}^{v}\}_{v=1}^{V}$ as,
$\bar{\mathbf{F}}^{v}=\bar{\mathbf{T}}^{v}\mathbf{F},$ (8)
@@
III-C Prior Knowledge-assisted Contrastive Learning
The similarities among samples defined from the original features are critical prior knowledge for feature representation learning [42]. In this section, we inject the prior knowledge into the view-common feature representation with useful self-supervised contrastive learning to further boost its representation capability. To this end, we formulate three key components, i.e., the clustering projection layer, mask-informed graph fusion, and contrastive loss.
Clustering projection layer: In this part, we first transfer the view-common representation into the cluster space with a single linear layer. Then, the QR decomposition [43] is introduced to guarantee the orthogonal property of the clustering indicator matrix $\mathbf{Y}\in\mathbb{R}^{N\times C}$, where $C$ is the number of clusters. And the clustering distributions of the incomplete multi-view data can be revealed in $\mathbf{Y}$.
Mask-informed graph fusion: To flexibly merge multiple similarity graphs constructed from IMvC data, view-specific graphs $\{\mathbf{A}^{v}\}_{v=1}^{V}$ are first transferred based on permutation matrices $\{\bar{\mathbf{T}}^{v}\}_{v=1}^{V}$ as,